OpenStudy (jmansedg):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A.2 B. -4/5 C. 5/4 D.-2

OpenStudy (anonymous):

Use formula m = (y2-y1)/ (x2-x1)

OpenStudy (anonymous):

where (x1, y1) = (1, -5) and (x2, y2) = (4, 1)

OpenStudy (anonymous):

m = ( 1- (-5)) / (4-1) m = (1+5) / (3) m= 6 / 3 m = 2 Answer